# Towards Graph Foundation Models: A Survey and Beyond

### v3

@@ 6.1.3 LLM-centric Methods
6.2 Pre-training
To train the model and enable it to handle both graph and text information, we need to train the model on a large amount of data. LLM and GNN can be pre-trained on textual data and graph data respectively. Furthermore, the GNN+LLM-based methods can further be pre-trained on data composed of graph and text. In this subsection, we category the pre-training strategies as GNN or LLM-based, and alignment-based.
6.2.1 GNN or LLM-based
Other frameworks leverage pre-trained LLMs to obtain text embeddings. The majority of existing models, such as GIANT [156], GraD [155], GraphFormer [167], GLEM [168], Text2Mol [173], MolCA [176], MoleculeSTM [174], and CLAMP [175] employ Masked Language Modeling (MLM) during pre-training. Some models, like TAPE and Graph-ToolFormer, opt for Language Modeling (LM) in the pre-training phase. Additionally, SimTeG integrates Text-Text Contrastive Learning (TTCL), a technique that leverages certain observed text pairs exhibiting more semantic similarity than randomly selected pairs during the pre-training phase as:
$\operatorname{Loss}_{\operatorname{TTCL}}=\mathbf{E}_{x,y^{+},y^{-}}\left[-\log\frac{\exp\left(k\left(x,y^{+}\right)\right)}{\exp\left(k\left(x,y^{+}\right)\right)+\exp\left(k\left(x,y^{-}\right)\right)}\right],$ (6)
6.2.2 Alignment-based
$+(1-y_{i})\log(1-k(\text{LLM}(\mu_{i}),\text{GNN}(\xi_{i}))),$
where $\mu_{i}$ is the text representation, $\xi_{i}$ is the graph representation, and $k$ is a score function to predict the activity of a molecule. The contrastive loss promotes the active molecules on a bioassay to have similar embeddings to the embedding of a specific bioassay, while ensuring that inactive molecules have dissimilar embeddings to it.
6.3 Adaptation
6.3.1 Fine-tuning
To adapt to the downstream tasks, some works [167, 162, 159, 156, 168, 160, 157, 167] utilize vanilla fine-tuning methods for node classification tasks. However, vanilla fine-tuning methods involve adjusting a broad range of model parameters, which can be computationally intensive and resource-demanding. So other works utilize parameter-efficient fine-tuning methods for downstream tasks, resulting in a more efficient and resource-friendly approach. Specifically, several studies [173, 174, 175] align the embedding space of GNN and LLM utilizing paired molecule graph-text data, while other research [155, 180, 172, 164, 161, 166] are tuned on TAGs with classification task. Additionally, some works works [178, 176, 181] adapt to downstream tasks by generating text captions or descriptions.
6.3.2 Prompt-Tuning
6.4 Discussion
Also, it is challenging to align LLMs and GNNs into a common representational space. To tackle this problem, it is essential to establish a robust standard for measuring the alignment between LLM and GNN representations. This standard should evaluate the degree to which the embeddings from both models capture similar semantic and structural information. Additionally, it is crucial to design effective methodologies for achieving this alignment. By doing so, we can ensure that the combined model leverages the strengths of both LLMs and GNNs, ultimately enhancing performance on various downstream tasks.
7 Challenges and Future Directions
7.1 Challenges about Data and Evaluation
7.1.1 Data Quantity and Quality
The improvements in data quantity and data quality are the key factors contributing to the effectiveness of foundation models [1]. At present, there is still a limited amount of open-source large-scale graph data [184, 185], and each dataset is primarily concentrated in a single domain. This poses a challenge to learn graph foundation models for diverse data domains. Hence, it is necessary to collect and organize a unified, massive dataset that covers graph data and related corpora across different domains. Additionally, if the graph data are noisy, incomplete, or not properly curated, it will negatively affect the performance of graph foundation models. To enhance the data quality of graph foundation models, efforts have been made to propose augmentation strategies from various perspectives, including graph structure learning, feature competion and label mixing, etc. However, since existing data augmentation techniques are typically tailored for individual GNN-based models, there is a need for further exploration on how to effectively augment graph data for LLM-based or GNN+LLM-based models.
7.1.2 Evaluation
With the help of natural language instructions and powerful generation capabilities, LLMs can support a variety of open-ended tasks [53]. This presents new opportunities for graph foundation models based on LLM. However, due to the lack of labels in open-ended tasks, evaluating the performance of GFMs in such tasks is a challenge. When using LLM as a language foundation model, the evaluation of its performance on open-ended tasks has evolved from human evaluation [61] to meta-evaluation [186]. The question of whether existing LLM evaluation methods [61, 186] can be applied to GFMs remains to be explored. Beyond evaluating the performance of graph foundation models, it is also worth evaluating their robustness, trustworthiness, or holistic performance, similar to the current practices for language foundation models [187, 188, 189].
7.2 Challenges about Model
7.2.1 Model Architectures
As mentioned above, the designs of backbone architectures and learning paradigms are crucial for the implementation of graph foundation models. Although this article has outlined some potential solutions, it does not rule out the possibility of better ones. For example, regarding the backbone architecture, recent works have proposed model architectures that go beyond the Transformer, offering improved performance [190] or interpretability [191]. However, it is still unknown whether these backbone architectures can be used for dealing with graph data. Additionally, when utilizing GNN+LLM-based models, it is worth exploring how to more effectively align the outputs of both models. Furthermore, there is limited research regarding the emergent abilities or neural scaling law [192, 193] of GNN-based [70] or LLM-based [71] graph foundation models. It is yet unclear whether GNN+LLM-based models may have greater potential for emergence. Given that current multimodal foundation models [194] primarily handle text, images, audio, and other modalities, it is an interesting research direction to explore whether GNNs can be employed to further expand the diversity of modalities covered by multimodal foundation models or enhance the capabilities of foundation models for multimodal learning [195].
7.2.2 Model Training
In order to achieve homogeneity and make effective use of pre-training data, it is crucial to design appropriate pretext tasks in pre-training. Unlike many language foundation models, which often use LM [61] or MLM [47] as pretext tasks, there are now various forms of pretext tasks tailored to different GFM model architectures. Whether each type of pretext task has its own applicable scope and whether there will be a unified pretext task is worth further exploration. After obtaining pre-trained graph foundation models, adapting them to downstream tasks is a vital concern. Apart from fine-tuning and prompting that are introduced in this article, there are other potential training techniques that can be applied to improve efficiency or update knowledge, such as knowledge distillation [196], reinforcement learning from human feedback (RLHF) [61] and model editing [197]. Whether the above-mentioned techniques can be applied to graph foundation models will be a focal point of future research.
7.3 Challenges about Applications
7.3.1 Killer Applications
In comparison to the outstanding performance of language foundation models in tasks like text translation [198] and text generation [199], whether graph foundation models can similarly catalyze groundbreaking applications in graph tasks is not yet clear. For scenarios that are well-suited for the application of GNNs, such as e-commerce [82] and finance [200], potential research directions include leveraging graph-based models integrated with LLMs to better support open-ended tasks [178], or enhancing the reasoning capabilities of LLMs through graph learning techniques [201]. Furthermore, graph foundation models have the potential to make breakthroughs in some emerging fields. For example, drug development is a time-consuming and costly process [202], and language foundation models have already been successfully used for related tasks like target identification and side effect prediction [1]. Given the 3D geometric structure of proteins [203, 204, 205], graph foundation models hold the promise of enhancing the drug discovery pipeline by leveraging their ability to model graph structure information [206], potentially speeding up the process further. Additionally, urban computing may also represent a crucial application scenario for graph foundation models. It is worth noting that traditional traffic prediction techniques have been primarily focused on addressing individual tasks such as travel demand prediction [207] and traffic flow prediction [208], lacking a comprehensive understanding of the entire transportation system. Given that the transportation system can be viewed as a spatio-temporal graph, graph foundation models hold the potential to capture the participation behavior of actors in the transportation system [209], thereby offering a unified approach to addressing various issues in urban computing.
7.3.2 Trustworthiness
Despite the strong performance of LLM-based foundation models, their black-box nature [210] introduces a host of safety concerns, such as hallucination and privacy leaks. The hallucination refers to the output appearing plausible but deviating from user input, context, or facts [211]. Existing research suggests that this phenomenon is associated with multiple factors, such as the model’s overconfidence in its own behavior [212] and the misunderstanding of false correlations [213]. Similarly, recent work has pointed out that pre-trained GNNs also pose certain trustworthy risks about fairness [214] and robustness against attacks [215]. Given the unique nature of graph data, we may require certain techniques to prevent or mitigate security risks on graph foundation models, such as confidence calibration [216] or counterfactual reasoning [217]. Additionally, given that existing research has indicated privacy risks in both GNN [218, 219] and LLM [220], enhancing the privacy of graph foundation models is also a critical concern. Some potential solutions include federated learning [221], RLHF [61] and red teaming [222], but whether these methods can be applied to graph foundation models is still unknown.
8 Conclusions

### v4

@@ A. Impact of Graph Data on GFMs
HiGPT [172] LLM-centric LM Parameter-Efficient FT
G-Retriever[171] LLM-centric LM Prompt-Tuning
TEA-GLM[170] LLM-centric LM Parameter-Efficient FT
HyperBERT[174] LLM-centric LM Parameter-Efficient FT
Graph Type. Based on the number of categories of nodes and edges in a graph, we can categorize graphs into homogeneous graphs and heterogeneous graphs. In homogeneous graphs, all nodes and edges belong to the same category. For example, in a social graph where nodes represent individuals (users) and edges represent friendship relationships, it is a homogeneous graph because all nodes are individuals and all edges represent friendship relationships. Heterogeneous graphs, on the other hand, have more than one type of nodes or edges, representing different types of entities and relationships [33]. For instance, an e-commerce graph may include nodes for users, products, and purchase relationships, forming a heterogeneous graph. For GFMs, handling heterogeneous graphs poses greater challenges and typically requires the design of specific backbone architectures and optimization objectives. Nonetheless, utilizing the meta-path based approach [223], a heterogeneous graph can be mapped into multiple homogeneous graphs, one for each meta-path. For example, one can apply the GFM trained on homogeneous graphs to each of these meta-path induced homogeneous graphs, separately, to get node embedding. Then, these embeddings on homogeneous graphs under different meta-paths can be fused together. However, beyond homogeneous graphs and heterogeneous graphs, there are some more complex types of graphs in the real world, such as dynamic graphs and hypergraphs [224], which poses additional challenges for GFM.
B. Impact of Graph Tasks on GFMs
Node-level Tasks. Node-level tasks refer to the classification, regression, or prediction performed on each node. Common node-level tasks include node classification, node regression, and clustering coefficient prediction. For example, in social networks, graph nodes can represent users, and node classification can be used to identify users from different social circles.
In summary, the format of tasks in graphs are highly diverse and can be categorized into three types: node-level, edge-level, and graph-level, each of which has wide-ranging applications. This undoubtedly increases the challenge of homogenization for GFMs. For example, in graph classification and node classification tasks on synthetic datasets, modeling structural information is often more crucial [228]. On the other hand, when dealing with node classification tasks on graphs with rich node features, modeling feature information becomes more important [228]. Furthermore, tasks that are more similar to each other will also have a lower transfer difficulty, implying that these tasks are more likely to be addressed using the same GFM. While increasing expressive power holds promise for improving the performance of many node-level, edge-level, and graph-level tasks [229], there is also some work suggesting that overly strong expressive power may not be necessary for graph generation tasks [230].
C. Details of approaches involved as GNN-based models
We categorize the GNN-based methods in Table I.
D. Details of approaches involved as LLM-based models
We categorize the LLM-based methods in Table II.
E. Details of approaches involved as GNN+LLM-based models
We categorize the GNN+LLM-based methods in Table III.